OpenStudy (anonymous):

Given the function g(x) = −2x + 4, find g(−5).

hero (hero):

g(x) = -2x + 4 g(-5) = -2(-5) + 4 g(-5) = 10 + 4 g(-5) = 14

OpenStudy (anonymous):

Woops, Hero is right

hero (hero):

That's why you should post all your steps